OpenStudy (anonymous):

Jack's pet dog has been prescribed x mg of a medicine per day. Each pill of the medicine has 10 mg of the medicine. Jack has to give his dog a total of 30 pills. If Jack gave his dog medicine for 9 days, use the expression below to interpret the situation. : 30 - 9 times x over 10 Part A: What does x over 10 represent? (3 points) Part B: What does the 9 times x over 10 represent? (3 points) Part C: What does the entire expression represent? (4 points)

OpenStudy (anonymous):

Part a: x/10 represents the amount of pills given divided by the amount of mg's in each pill. Part b: 9*x/10 represents the number of pills given in each day of the nine days and how many mg' are in the total dosage for the nine days. Part c: If Jack has to give his dog 30 pills in nine days the entire expression represents the number of pills given to the dog each day and how many mg's of medicine the dog is consuming.

OpenStudy (anonymous):

thanks

OpenStudy (gorv):

total pills = 30 one day dos = x/10 9 day doze= 9*x/10 which shpould be equal to 30 30=9x/10